### v1

@@ Abstract
Keywords:
Knowledge Tracing Educational Data Mining Student Simulator
@@ 2 Problem Formulation
In contrast to existing KT models, we focus on a multi-step prediction paradigm, which is more in line with real-world applications. Specifically, this approach involves predicting the student’s response $r_{T}$ using previously predicted responses $[\hat{r}_{t}]_{t=1}^{T-1}$, which can be expressed as:
@@ 3.2 Generator
$\mathbf{Q}=\mathbf{v}_{q_{T}},\mathbf{K}$ =$=\{\mathbf{v}_{q_{1}},\cdots,\mathbf{v}_{q_{T-1}}\},\mathbf{V}=\{\mathbf{h}_{1},\cdots,\mathbf{h}_{T-1}\},$
@@ 3.3 Discriminator
The discriminator takes the question sequence $[q_{i}]_{i=1}^{T}$, response sequence $[r_{i}]_{i=1}^{T}$, and positions $[o_{i}]_{i=1}^{T}$ as inputs. We define the question-position matrix $\mathbf{v}_{q}$ and response-position matrix $\mathbf{v}_{r}$ as:
@@ 3.4.3 Adversarial Learning
The general training algorithm is illustrated in Appendix LABEL:alg:AdvKT. We employ an alternating update strategy to iteratively update the discriminator and the generator. When updating the discriminator, we first generate four types of training data with rules, and compute the loss function in Eq. (19), through which the augmented discriminator learns gain robust capabilities. During the training of the generator, it receives rewards from the discriminator, guiding it to simulate more realistic responses, while also being trained by the additional losses in Eq. (15).
@@ 4.1.3 Implementation Detail
The maximum length of each student’s learning sequence is $200$. The dimension of the hidden state of GRU is set to $64$. The question embedding, concept embedding, response embedding and position embedding all have a dimension of $64$. We utilize a $4$-headed Transformer in the discriminator. The value of $\gamma$ in Eq. (12) is chosen in {$0.9$, $0.93$, $0.95$ , $0.98$}. In Eq. (15), $\lambda_{1}$ is set to $1000$, and $\lambda_{2}$ is chosen from $\{0,1\}.$ The optimizer used is Adam [adam]. We update the discriminator every $2$ epochs. The learning rate of the generator is $0.001$, while the learning rate of the discriminator is $0.005$. The memory overhead incurred during training by AdvKT is comparable to that of recent knowledge tracing models, with AdvKT requiring approximately 15,000 MiB, while LBKT [32] utilizes around 12,600 MiB. During inference stage, AdvKT only requires the use of the generator, thereby reducing GPU memory requirements by half. Our model is implemented on PyTorch and is available for reviewers  [^2] Source code for AdvKT: https://anonymous.4open.science/r/AdvKT-3E60/ and will be publicly available upon the acceptance of this work.
Table 2: Overall performance of AdvKT on four public datasets. The best performing and second-best models are denoted in bold and underlined. indicates p-value $<0.05$ in the significance test.
@@ 4.4 Mitigation of Data Sparsity
In Figure 4, we visualize the distribution of $(q,r)$ pairs in (i) original data $\mathcal{R}$ in the dataset, (ii) augmented data $\mathcal{T}$ and (iii) generative data $\mathcal{E}$ on four datasets. Our discriminator encodes the entire sequence and obtains hidden state representations. These representation vectors are reduced to two dimensions using t-SNE [21].
@@ 4.6 Insight of Adversarial Learning
Generative data $\mathcal{G}$ plays a crucial role in training the generator alongside the real dataset. The quality and diversity of these synthetic sequences directly influence the supervision signal ($\mathcal{L}_{\text{adv}}$) obtained by the generator. that the generator receives. In Figure 6, we present a generated synthetic sequence along with its most similar learning sequences from the dataset. Four subsequences are found within the original sequences, highlighting a certain similarity between the generated and original sequences. However, each generated sequence remains distinct from any single original sequence.
@@ 5.1 Knowledge Tracing
To investigate students’ learning patterns, numerous knowledge tracing models have been proposed under a single-step training paradigm.
Deep Knowledge Tracing (DKT) [26] pioneers using the deep learning-based knowledge tracing methods. It uses the hidden states of long short-term memory (LSTM) [13] networks to describe the students’ knowledge states. After DKT, DHKT [31] and other variants of DKT introduce extra information like question-skill relations and additional regularization terms to improve the performance. Inspired by Key-Value Memory Network (KVMN) [22], several works [34, 1] use key-matrix and value-matrix to represent concepts and student’s mastery level. GKT randomly builds a similarity graph of skills, GIKT [33] applies the graph of skills and questions to obtain better representations of questions. After the transformer architecture was proposed, some works [4, 29] tried to utilize the attention mechanism in KT tasks.

### v2

@@ 2 Problem Formulation
In contrast to existing KT models, we focus on a multi-step prediction paradigm, which is more in line with real-world applications. Specifically, this approach involves predicting the student’s response $r_{T}$ using previously predicted responses $\hat{r}_{t}]_{t=1}^{T-1}$, which can be expressed as:
@@ 3.2 Generator
$\mathbf{Q}=\mathbf{v}_{q_{T}},\mathbf{K}$ $=\{\mathbf{v}_{q_{1}},\cdots,\mathbf{v}_{q_{T-1}}\},\mathbf{V}=\{\mathbf{h}_{1},\cdots,\mathbf{h}_{T-1}\},$
@@ 3.3 Discriminator
The discriminator takes the question sequence $q_{i}]_{i=1}^{T}$, response sequence $r_{i}]_{i=1}^{T}$, and positions $o_{i}]_{i=1}^{T}$ as inputs. We define the question-position matrix $\mathbf{v}_{q}$ and response-position matrix $\mathbf{v}_{r}$ as:
@@ 3.4.3 Adversarial Learning
The general training algorithm is illustrated in Appendix 0.A. We employ an alternating update strategy to iteratively update the discriminator and the generator. When updating the discriminator, we first generate four types of training data with rules, and compute the loss function in Eq. (19), through which the augmented discriminator learns gain robust capabilities. During the training of the generator, it receives rewards from the discriminator, guiding it to simulate more realistic responses, while also being trained by the additional losses in Eq. (15).
@@ 4.1.3 Implementation Detail
The maximum length of each student’s learning sequence is $200$. The dimension of the hidden state of GRU is set to $64$. The question embedding, concept embedding, response embedding and position embedding all have a dimension of $64$. We utilize a $4$-headed Transformer in the discriminator. The value of $\gamma$ in Eq. (12) is chosen in {$0.9$, $0.93$, $0.95$ , 0.980.98}. In Eq. (15), $\lambda_{1}$ is set to $1000$, and $\lambda_{2}$ is chosen from $\{0,1\}.$ The optimizer used is Adam [1]. We update the discriminator every $2$ epochs. The learning rate of the generator is $0.001$, while the learning rate of the discriminator is $0.005$. The memory overhead incurred during training by AdvKT is comparable to that of recent knowledge tracing models, with AdvKT requiring approximately 15,000 MiB, while LBKT [33] utilizes around 12,600 MiB. During inference stage, AdvKT only requires the use of the generator, thereby reducing GPU memory requirements by half. Our model is implemented on PyTorch and is available on Github [^2] Source code for AdvKT: https://github.com/fulingyue/AdvKT.
Table 2: Overall performance of AdvKT on four public datasets. The best performing and second-best models are denoted in bold and underlined. indicates p-value <0.05<0.05 in the significance test.
@@ 4.4 Mitigation of Data Sparsity
In Figure 4, we visualize the distribution of $(q,r)$ pairs in (i) original data $\mathcal{R}$ in the dataset, (ii) augmented data $\mathcal{T}$ and (iii) generative data $\mathcal{E}$ on ASSIST09. Visualization of distribution on other datasets are demonstrated in Appendix B. Our discriminator encodes the entire sequence and obtains hidden state representations. These representation vectors are reduced to two dimensions using t-SNE [22].
@@ 4.6 Insight of Adversarial Learning
Generative data $\mathcal{G}$ plays a crucial role in training the generator alongside the real dataset. The quality and diversity of these synthetic sequences directly influence the supervision signal ($\mathcal{L}_{\text{adv}}$) obtained by the generator. In Figure 6, we present a generated synthetic sequence along with its most similar learning sequences from the dataset. Four subsequences are found within the original sequences, highlighting a certain similarity between the generated and original sequences. However, each generated sequence remains distinct from any single original sequence.
@@ 5.1 Knowledge Tracing
To investigate students’ learning patterns, numerous knowledge tracing models have been proposed under a single-step training paradigm. Deep Knowledge Tracing (DKT) [27] pioneers using the deep learning-based knowledge tracing methods. It uses the hidden states of long short-term memory (LSTM) [14] networks to describe the students’ knowledge states. After DKT, DHKT [32] and other variants of DKT introduce extra information like question-skill relations and additional regularization terms to improve the performance. Inspired by Key-Value Memory Network (KVMN) [23], several works [35, 2] use key-matrix and value-matrix to represent concepts and student’s mastery level. GKT randomly builds a similarity graph of skills, GIKT [34] applies the graph of skills and questions to obtain better representations of questions. After the transformer architecture was proposed, some works [5, 30] tried to utilize the attention mechanism in KT tasks.
@@ 6 Conclusion
{credits}
6.0.1 Acknowledgements
The Shanghai Jiao Tong University team is partially supported by National Key R&D Program of China (2022ZD0114804), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and National Natural Science Foundation of China (62322603, 62177033).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Adversarial Training Algorithm
We present an adversarial training paradigm to train the generator and discriminator together. The overall training algorithm is illustrated in Algorithm 1. We alternatingly update the discriminator and the generator until convergence.
1
2Initialize the Generator $\mathcal{G}$, Discriminator $\mathcal{D}$ randomly;
3 while not converged do
4
5 foreach batch data $\mathcal{R}$ do
6    if Update Discriminator then
7         /* Update Discriminator */
8         Generate augmented data $\mathcal{T}$;
9         Generate generative data $\mathcal{E}$ with generator $\mathcal{G}$;
10         Generate reversed label data $V$;
11         Compute the loss function Eq. (19) and update $\mathcal{D}$;
12
13     /* Update Generator */
14      Compute the supervise learning loss Eqs. (11) and (14);
15      Generate generative data $\mathcal{E}$;
16      Compute the policy gradient loss Eq. (12) using $\mathcal{E}$;
17      Update $\mathcal{G}$ with Eq. (15);
18   end foreach
19
20 end while
Algorithm 1 Adversarial Training of AdvKT
EdNet
Slepemapy
Junyi
Figure 7: Distribution of question-response pair visualized with two-dimensional t-SNE. Green points are real data in dataset. Blue points correspond to data generated by popular data augmentation methods (mask, crop, permute, replace). Red points represent generative data. The three lines of images are visualizations of EdNet, Slepemapy and Junyi.
Appendix 0.B Data Distribution Visualization
In Figure 7, we visualize the distribution of $(q,r)$ pairs in (i) original data $\mathcal{R}$ in the dataset, (ii) augmented data $\mathcal{T}$ and (iii) generative data $\mathcal{E}$ on EdNet, Slepemapy and Junyi. The figure illustrates that even though different datasets exhibit significant variations in data distribution, real data consistently suffer from the issues of data sparsity and non-uniformity. In EdNet’s real data, there are three noticeable gaps, and the points along the data boundaries are also unevenly distributed. As for Slepemapy’s real data, while the boundaries appear relatively uniform, there exists a distinct hole within. Junyi’s data, on the other hand, skews towards the top left corner, with sparse points in the bottom right corner. Our four positive sample augmentation methods first expand the dataset with real data, ensuring that the augmented data still aligns with the real data. Additionally, generative data can fill in the gaps in real data, resulting in a more evenly distributed training dataset across the entire space. Therefore, our data augmentation on the discriminator’s training data could well solve the data unevenness and sparsity.